Two hockey players push off from a clinch, recoiling from rest. Jean-Claude has an inertia that is 60% greater than Pierre's inertia. After they push off and move across the ice in opposite directions, they lose kinetic energy at the same rate until they come to a stop.

1.What is the ratio of the distances which players traveled?
1
Expert's answer
2020-03-10T11:27:07-0400
"m_2=1.6m_1"

From the conservation of momentum:


"m_1v_1=m_2v_2\to v_1=1.6v_2"

"K_{10}=0.5m_1v_1^2"

"K_{20}=0.5m_2v_2^2=\frac{K_{10}}{1.6}=0.625K_{10}"

"\frac{d_{JC}}{d_P}=\frac{K_{20}}{K_{10}}=0.625"